the first resonant length of a resonance tube is 40cm and the second resonant length is 122cm. find the third resonant length of resonance tube?
1
Expert's answer
2017-10-11T15:37:07-0400
The first resonant length is
L_1=λ/4
The second resonant length is
L_3=3 λ/4
The third resonant length of resonance tube is
L_5=5 λ/4=5/3 L_3=5/3 (122cm)=203 cm.